Assume the NPV of a project is $1.5 million. The project is expected to last five years. What is the equivalent annual benefit if the discount rate is 8%?
Answer #1

NPV =1,500,000
Number of Years =5
Discount Rate =8%
Equivalent Annual Cost =NPV/(1-(1+r)^-n)/r) =1500000/(1-(1+8%)^-5)/8%) =375,684.68